Finding the slope of the line through each pair of points
(19,-16), (-7,-15)

Answer by CharlesG2(834)

Finding the slope of the line through each pair of points
(19,-16), (-7,-15)
slope = m = rise/run = (y2-y1)/(x2-x1)
m = (-15 - -16)/(-7 - 19)
m = (-15 + 16)/(-26)
m = 1/(-26)
m = -1/26